1 In a graduate teacher college, a survey was conducted to determine the proportion of students who want to

major in Mathematics. If 378 out of 900 students said Yes, with 95% confidence, what interpretation can we

make regarding the probability that all students in the teacher graduate college want to major in Mathematics.


1
Expert's answer
2022-05-30T16:24:17-0400

The sample proportion is computed as follows, based on the sample size "n = 900" and the number of favorable cases "X = 378:"


"\hat{p}=\dfrac{X}{n}=\dfrac{378}{900}=0.42"


The critical value for "\alpha = 0.05"  is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI(proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.42-1.96\sqrt{\dfrac{0.42(1-0.42)}{900}},"

"0.42+1.96\sqrt{\dfrac{0.42(1-0.42)}{900}})"

"=(0.388,0.452)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.388 < p < 0.452," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.388, 0.452)."